Complete the ordered pair so that it satisfies the given equation.
3x – 5y = –18: (___, 3)

Answer by nerdybill(7384)

Your equation:
3x – 5y = –18
.
Your "ordered pair":
(___, 3)
which is in the order of (x, y)
This means, they want to know what 'x' is when y=3
.
Simply substitute the value of y into your equation and solve for 'x':
3x – 5y = –18
3x – 5(3) = –18
3x – 15 = –18
3x = –3
x = -1
.
Therefore, the "ordered pair" is:
(-1, 3)